### v1

# TikTokActions: A TikTok-Derived Video Dataset for Human Action Recognition

Abstract
The increasing variety and quantity of tagged multimedia content on platforms such as TikTok provides an opportunity to advance computer vision modeling. We have curated a distinctive dataset of 283,582 unique video clips categorized under 386 hashtags relating to modern human actions. We release this dataset as a valuable resource for building domain-specific foundation models for human movement modeling tasks such as action recognition. To validate this dataset, which we name TikTokActions, we perform two sets of experiments. First, we pretrain the state-of-the-art VideoMAEv2 with a ViT-base backbone on TikTokActions subset, and then fine-tune and evaluate on popular datasets such as UCF101 and the HMDB51. We find that the performance of the model pre-trained using our TikTok dataset is comparable to models trained on larger action recognition datasets(95.3% on UCF101 and 53.24% on HMDB51). Furthermore, our investigation into the relationship between pre-training dataset size and fine-tuning performance reveals that beyond a certain threshold, the incremental benefit of larger training sets diminishes. This work introduces a useful TikTok video dataset that is available for public use and provides insights into the marginal benefit of increasing pre-training dataset sizes for video-based foundation models.
1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as surveillance, interactive gaming, and healthcare. In the era of social media platforms, TikTok has emerged as a significant repository of human action videos. We have compiled a dataset of 283,582 TikTok video clips that encompass a range of human activities, including dance, sports, fitness, hand gestures, and daily routines. This collection aims to represent various human actions in a manner that is relevant to real-world scenarios and cultural contexts. Importantly, this dataset is uniquely curated to include well-defined but non-standard human actions, filling gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models. We developed a model for human action recognition trained using the curated TikTok video clips. Pre-training on a dataset with varied and authentic content is intended to more accurately reflect the spectrum of naturalistic human actions, aiming to address the limitations of conventional datasets by incorporating elements of contemporary behavior. In our initial evaluations, presented later in this paper, we find that utilizing our TikTok dataset may enhance the development of targeted human action recognition models in the field of computer vision.
Moreover, the nature of our data source, TikTok, presents another advantage. Kinetics[4] derives its content from YouTube, featuring a combination of professional and amateur videos, whereas TikTok’s database consists primarily of user-generated material. This means that our dataset is inherently rich with real-world scenarios: from spontaneous camera movements and varied illumination conditions to diverse backgrounds and authentic human interactions. This rawness provides a genuine challenge and opportunity for training robust models capable of understanding human actions in a plethora of real-world conditions.
We conducted a series of experiments to assess the quality of our dataset and its potential to serve as a basis for pre-training foundation models for action-related tasks. We used a well-established state-of-the-art model VideoMAE V2[9], and pre-trained it on our TikTokActions dataset. We then fine-tuned the model on two well-known video action recognition datasets: UCF101[7] and HMDB51[5]. Our experiments provided valuable insights: the performance achieved was comparable to models that were both pre-trained and fine-tuned solely on UCF101 and HMDB51. This underscores the utility of the TikTokActions dataset to serve as a useful pre-training dataset for action-related tasks.
We next delved into understanding the relationship between the quantity of pre-training videos and the performance on the downstream action recognition task. By sampling subsets of our TikTok dataset, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, pre-training VideoMAE V2[9] on the sampled subset, and subsequently fine-tuning each on the benchmark datasets, we observed that as long as the number of pre-training videos remains above a certain threshold, the performance remains comparable to massive pre-training datasets. This provides a valuable lesson when pre-training foundation models: the benefits of increasing the pre-training dataset size diminish as the dataset increases.
2 Related Work
2.1 Human Action Recognition Datasets
The evolution of action recognition research has been strongly influenced by the availability of benchmark datasets. The UCF101[7] dataset, with its diverse collection of human actions, has been a cornerstone for early research in this domain, guiding and benchmarking numerous computer vision innovations applied to action recognition. Similarly, the HMDB-51[5] dataset, despite its limited size, has been pivotal in providing a varied set of human actions captured under realistic conditions.
The Kinetics Human Action Video dataset[4], introduced in 2017, marked a significant milestone in this field. With its vast collection of YouTube-sourced videos encompassing 400 distinct human action classes, Kinetics has become an indispensable resource for researchers, setting a higher benchmark and promoting advanced research in action recognition.
Another noteworthy dataset is the Something-Something[3] collection. Its unique focus on human-object interactions captured in real-world scenarios has offered researchers a different perspective, emphasizing the complexity and diversity of everyday actions.
These datasets, along with several others, have collectively laid the foundation for the current advancements in action recognition, facilitating the development and evaluation of state-of-the-art models and techniques. Recognizing the rapid evolution of content on platforms like TikTok and the emergence of diverse, contemporary actions and scenarios, we are motivated to develop a comprehensive dataset highlighting the latest trends and user behaviors specific to new short-video social media platforms. Our aim was to provide the community with a dataset with well-defined yet non-standard human actions that complement common actions already represented in current action recognition datasets, thus adding variety to the actions that action recognition models are currently evaluated for.
2.2 Computer Vision Foundation Models
Foundation models have become essential tools in the field of computer vision. The advent of models like Florence[10] showcases the versatility and scope of these models. Specifically, Florence encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences. Additionally, its effectiveness spans various modalities, from RGB imagery to textual data, establishing new standards in tasks like classification, object detection, retrieval, and particularly in action recognition.
Another significant model in this domain is VideoMAE V2[9]. Introduced at CVPR 2023, this self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos. With its innovative dual masking strategy and progressive training approach, VideoMAE V2 sets new standards in action recognition. Empirical evaluations demonstrate its superior performance across multiple benchmarks, emphasizing its effectiveness in real-world scenarios. We anticipate that our extensive TikTok dataset will serve as a valuable resource for the continued development and refinement of computer vision foundation models, contributing to the advancement of this rapidly evolving field.
3 Dataset Construction
3.1 Dataset Overview
We introduce the TikTokActions dataset, designed for human action recognition. Derived from TikTok videos, this dataset comprises over 280,000 video clips from 386 hashtags that capture a wide spectrum of human actions in various scenarios. The dataset spans diverse categories, from dance movements and fitness exercises to hand gestures and daily activities. The aim of TikTokActions is to provide a robust and comprehensive foundation for training and evaluating human action recognition models in realistic settings.
3.2 How the Dataset was Built
To identify a list of hashtags to mine on TikTok, we initially consulted several commonly used datasets related to human actions, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4]. After conducting a thorough comparison of these datasets, we identified four categories of videos that can best represent various human actions: dance, sports, fitness, and kinetics. Following these categorizations, we proceeded to identify several hashtags under each category. During the process of structuring our hashtag list, we drew insights from the hashtags commonly employed by TikTok users on the platform. This comparative analysis yielded six fundamental principles that guided our approach to selecting hashtags to mine for data: generalization, specification, noun substitution, noun-verb order, hashtags consolidation, and social/cultural consideration. The last principle enabled our team to include the fifth category unique to social media, encompassing a wide variety of challenge types of actions originating from distinct social and cultural contexts. Additionally, beyond these six principles, we placed significant emphasis on the number of views under each hashtag, recognizing that popularity and engagement levels are crucial indicators of the relevance and impact of the content for each hashtag. We set a 5 million view minimum threshold for hashtag inclusion, reflecting a focus on content demonstrating substantial public engagement and visibility, indicative of broad influence.
3.2.1 Generalization/Using the umbrella terms
When comparing the hashtags we generated with those used by TikTok users, we observed that hashtags related to specific actions can often be represented in a very broad or inclusive manner on the platform. For instance, the hashtag #basketball can serve as an umbrella term encompassing a wide array of hashtags, such as #playingbasketball, #dunking, #dribbling.
3.2.2 Specification
There are situations where specificity is crucial. For instance, within the broader classification of Ballroom Dance, distinct dance styles such as Waltz, Quickstep, and Tango exist. Instead of solely using the hashtag #ballroomdance, we include specific hashtags, including #chacha, #salsa, #tango, and #waltz, to ensure our datasets encompass pertinent human actions inherent to distinct dance styles.
3.2.3 Noun Substitution
While gathering short TikTok videos, we observed that a more effective approach for identifying the desired content is to employ a single noun instead of an entire verb phrase. For instance, the use of the hashtag #piano suffices to compile a more extensive collection(see Table  1) of short videos instead of using #pianoplaying.
Hashtags Number of Views
#piano 44.4 Billion
#playingpiano 29.4 Million
Table 1: Table of Hashtags and Their Corresponding Number of Views
Figure 1: Flowchart depicting the hashtag selection and modification process.
3.2.4 Noun-Verb Order
We noticed a preference among users for using noun phrases over verb phrases when describing their activities. For example, TikTok users frequently use hashtags like #shoerepair instead of verb phrases like #repairingshoes.
3.2.5 Hashtag Consolidation
When composing the hashtag list, we streamlined similar hashtags into a single one if they denote the same type of human action, prioritizing those with higher views on TikTok. For example, we retain #cuttingcakes and remove similar hashtags like #cutfruit, #applecutting, and #breadcutting.
3.2.6 Social/cultural Considerations
Culturally and socially relevant hashtags were included when crafting the hashtag list. For instance, the ”flower challenge” originated as a dance challenge on TikTok initiated by a K-pop star, and TikTok users subsequently adopted the hashtag #flowerchallenge when sharing their dance videos.
In Figure 1, we show an example flowchart of the hashtag selection and modification process. Our final and complete list of hashtags along with video numbers is in Appendix A.
3.3 Data Collection and Processing
We constructed the TikTokActions dataset using the official TikTok research API. For each hashtag, approximately 900 videos and corresponding metadata were mined. Each video underwent processing with tools like PySceneDetect and YOLO v8[6] producing a JSON file for every video, thereby accumulating metadata for 340,000 videos. The metadata encompassed various attributes, including the title, duration, and scene distribution based on PySceneDetect. We describe details about the preprocessing procedure below.
3.3.1 PySceneDetect and YOLOv8 Preprocessing
Our pre-processing approach began with the use of PySceneDetect to analyze the scene composition of each video. PySceneDetect was employed to identify distinct scenes within the videos, enabling us to focus on the longest scene clips. The rationale behind selecting the longest scenes was our hypothesis that these segments have the highest probability of containing human presence.
Once the longest scenes were extracted, we then utilized the YOLOv8[6] model to detect human figures in these video segments. The primary goal of this method was to accurately identify videos containing human subjects. The procedure comprised several distinct phases, each fulfilling a specific aspect of the preprocessing objective.
In the frame extraction phase, the total frame count of each video clip was determined using the ffprobe tool. Given the considerable frame count in many videos, it was computationally impractical to process each frame. Consequently, we resorted to subsampling. Typically, a fixed number of frames, defaulting to 10, were extracted at regular intervals. For a video containing 101 frames, every 10th frame was extracted starting from the 10th frame.
The next step was human detection. Here, the extracted frames underwent inference using the YOLOv8[6] model, initialized based on Ultralytics’ YOLO framework. The primary task was to review the detections and identify frames with human presence. Frames with at least one ”person” label contributed to the cumulative count of frames with human subjects for that particular video.
Lastly, the post-processing steps included thresholding. For a video to qualify as containing humans, it was required to surpass a pre-defined frame threshold with human detections, set to 4 by default. Videos below this threshold were earmarked for subsequent human scrutiny. Following the analysis of all videos, a results compilation phase calculated aggregated statistics, providing a clear picture of the number of videos with humans. Additionally, any extracted frames were systematically removed post-inference to maintain a clean working environment. In Figure  2, we show the process of the generation of clips for each video.
Figure 2: Flowchart depicting the clip extraction process from raw videos mined on TikTok.
4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is crucial to understanding the efficacy of a model. In this section, we evaluate the performance of our model, which was initially pre-trained on VideoMAE V2[9] with TikTokActions subset and subsequently fine-tuned on two datasets: UCF-101[7] and HMDB51[5].
4.1 Pre-training on VideoMAE V2
VideoMAE V2[9] represents a state-of-the-art masked autoencoding framework with a billion-level ViT backbone. This framework has been instrumental in enhancing performance across a diverse range of video-related tasks, including video action recognition, action detection, and temporal action detection. The generalization ability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been verified across diverse downstream tasks, showcasing its robustness and adaptability.
Given the strength of VideoMAE V2[9], we utilized it for our pre-training phase. Table 2 displays the hyperparameters used for this phase. The weights of VideoMAE V2 were randomly initialized.
Table 2: Hyperparameters for Pre-training on VideoMAE V2
Parameter Value
Mask Ratio 0.9
Decoder Mask Ratio 0.5
Model Vit_base
Decoder Depth 4
Batch Size 32
Number of Frames 16
Sampling Rate 4
Number of Samples 2
Number of Workers 24
Optimizer adamw
Learning Rate 1e-3
Optimizer Betas 0.9, 0.95
Warmup Epochs 20
Total Epochs 200
We conducted our training on a single NVIDIA A100 40GB GPU. This constraint necessitated a reduction in the scale of our dataset; consequently, we focused on a randomly selected subset comprising 10,000 videos, selected from the total collection of 280,000. This approach allowed us to manage the computational demands while still providing a representative sample for training and evaluating the model’s performance. This constraint also enabled us to study the utility of self-supervised pre-training on smaller datasets and compare it to prior work which are pre-trained on larger datasets.
4.2 Fine-Tuning on Established Datasets
Parameter Value
Batch Size 8
Input Size 224
Short Side Size 224
Number of Frames 16
Sampling Rate 4
Number of Workers 24
Optimizer adamw
Learning Rate 1e-3
Drop Path 0.35
Layer Decay 0.9
Optimizer Betas 0.9, 0.999
Weight Decay 0.05
Total Epochs 100
Test Number of Segments 5
Test Number of Crops 3
Table 3: Hyperparameters for Fine-tuning
Dataset Top-1 (%) Top-5 (%)
HMDB-51[5] 53.24 89.73
UCF-101[7] 95.3 99.54
Table 4: Fine-tuning Results With 10,000 Tiktok Actions videos Pre-training
After the initial pre-training on VideoMAE V2[9], the model was fine-tuned on established datasets. The hyperparameters for the fine-tuning phase are provided in Table 3. The results of the fine-tuning on various datasets, UCF-101[7] and HMDB51[5] are consolidated in Table 4. It showcases the effectiveness of the TikTok Actions dataset in model pre-training. Notably, the model achieved a high Top-1 accuracy of 95.3% on UCF-101 and 53.24% on HMDB-51, demonstrating its strong generalization capabilities across diverse action recognition tasks.
5 Extrapolating the Required Pre-Training Dataset Size
In Figure LABEL:fig:accplot, we find a monotone increasing accuracy score versus data volume. In the spirit of Floares et al. [2], we attempt to extrapolate the number of preprocessed videos for a desired accuracy by curve fitting to the number of data points trained so far. Because the accuracy maxes out at 100%, we choose to fit curves that are monotonically increasing with a horizontal asymptote at $11$. This makes any CDF of a random variable a viable choice for curve fitting. By iterating over the random variables in Scipy [8], we choose the top 10 distributions that fit the accuracy curve. Table 5 displays the number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
We validated these projections by pre-training a model with 9157 data points and fine-tuning the model using the same methods applied in the initial projections. The outcome of this validation was notably consistent with our predictive model. The model achieved an accuracy of 94.93%, closely aligning with the expected performance as indicated by the inverse CDF projections for this dataset size.
This validation not only reinforces the reliability of our statistical model but also highlights the practicality of using such predictive techniques in determining dataset size requirements. The fact that a dataset size of just 9157 points could yield an accuracy so close to 95% – a benchmark often targeted in action recognition tasks – is a testament to the efficiency of our approach. These results suggest that the requirement for massive datasets in the pre-training phase may be reconsidered, especially when computational resources are limited. This could mark a significant shift in how datasets are compiled and utilized in machine learning, particularly in fields where data acquisition is a challenge.
Table 5: Number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
Model Desired Accuracy
95% 96% 97% 98% 99%
t
9157 12388 18323 31892 82574
Burr XII
9090 12209 17857 30511 76240
InvGamma
9072 12166 17750 30204 74859
BetaPrime
9072 12165 17748 30200 74843
F
9072 12166 17749 30203 74855
GenPareto
9072 12164 17744 30184 74759
Lomax
9072 12164 17744 30184 74759
LogLaplace
9072 12164 17744 30184 74760
Pareto
9072 12164 17743 30183 74755
InvWeibull
9043 12101 17596 29772 72915
6 Discussion
The development of the TikTok-sourced dataset aimed to contribute to the action recognition field by including a variety of well-defined but non-standard human actions. These actions were selectively curated to complement the actions typically found in existing datasets, rather than duplicating them. This approach in dataset composition was chosen to diversify the range of human actions available for training and evaluating action recognition models.
By integrating these unique and diverse actions from the TikTok platform, our dataset broadens the scope of actions recognizable by machine learning models. This expansion is not merely a quantitative increase in data but a qualitative enrichment that challenges and enhances the adaptability and comprehensiveness of action recognition algorithms. This contribution is particularly significant in light of the evolving nature of human actions in digital media, where new and unconventional actions frequently emerge.
We also made auxiliary findings with notable implications: we observed a diminished necessity for large pre-training dataset sizes. Even with a relatively smaller subset of the TikTokActions dataset, the model performance remains competitive. This challenges the prevailing notion in action recognition research that larger datasets are invariably superior for pre-training. The implications of this finding indicate a potential shift in the paradigm of dataset compilation and model training. It points towards the efficiency of a smaller, well-curated dataset that can significantly reduce computational resources while maintaining high accuracy, which may be preferred over a larger but less valuable dataset.
One of the primary limitations of this study is the incomplete pre-training on the full TikTokActions dataset. Due to computational constraints, our experiments were conducted on a subset of 10,000 videos, which might not fully represent the diversity and complexity of the full dataset. Additionally, while the dataset performed well on benchmarks like UCF101[9] and HMDB51[5], its efficacy across other popular action recognition datasets remains untested.
Another significant limitation is the absence of ground truth labels for each clip in the TikTokActions dataset. Although the videos were curated scientifically, the lack of verified labels poses a challenge for more rigorous supervised learning approaches, potentially impacting the precision of the training and evaluation processes.
Future research endeavors can utilize the entirety of or subsets of the TikTokActions dataset for pre-training, aiming to enhance model performance and generalization across action recognition tasks. Another opportunity is the integration of TikTokActions with other datasets such as Kinetics[4], which will provide insights into the unique features and potential synergies of these datasets. This comparison is crucial for validating TikTokActions’ effectiveness and identifying areas for expansion or enhancement.
An emerging area of exploration is the application of weekly self-supervised learning techniques. These methods, which involve regular, unsupervised updates to the model based on new data influxes, could significantly enhance the adaptability and accuracy of the models trained on TikTokActions.
Furthermore, employing crowd-sourcing to acquire ground truth labels for the TikTokActions dataset presents a promising opportunity. This initiative would enable researchers to conduct more supervised learning studies, potentially leading to more accurate and reliable models. By combining crowd-sourced annotations with sophisticated learning algorithms, the field can expect to see substantial advancements in action recognition capabilities and applications.
Appendix A Appendix
In the appendix, we will present additional insights into our dataset, specifically focusing on visual and statistical aspects. This includes select screenshots of our video clips, which exemplify the diversity and context of the actions captured within our dataset. Alongside these visuals, we will include the figure, which provides the number distributions of videos for each hashtag. Furthermore, we will present the distribution of video durations, offering a comprehensive view of the temporal characteristics of our clips. These elements combined will give a more complete picture of the dataset’s richness.
Appendix B Sample Clips
In Figure 3, we present the extensive diversity of the TikTokActions dataset, which captures a wide range of human activities tagged across multiple domains such as fitness, sports, and dance. The collection features a variety of video types that include full-body human actions, clips focusing on hand movements, and scenarios involving multiple individuals. The presence of such varied content provides a rich learning ground for self-supervised learning models to extract and learn a comprehensive set of features from human actions, gestures, and interactions, aiding in the enhancement of action recognition capabilities.
Figure 3: Example classes from the TikTokActions dataset.
Appendix C Dataset Composition
Our dataset offers a diverse array of video counts across different hashtags. Figure 4 illustrates the distribution of videos across various hashtags, indicating the prevalence of certain actions within our dataset. On average, each hashtag is associated with approximately 735 videos. The ’sax’ category stands out with the maximum count of 938 videos, while the ’mopping’ category has the minimum, with 325 videos.
Figure 4: Bar chart showing the number of videos per hashtag in our dataset, demonstrating the frequency and distribution of action categories.
Conversely, Figure 5 displays a pie chart that categorizes the dataset into four video count ranges: 200-400, 400-600, 600-800, and 800+ videos per category. It is noteworthy that a significant portion, 53.4%, of the hashtags have a video count that falls within the 600-800 range.
Figure 5: Pie chart categorization of video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.
Appendix D Dataset Duration
In the composition of the TikTokActions dataset, a key focus was on ensuring that each video clip is sufficiently long to capture a complete action, while also maintaining a manageable length for analysis. To this end, we established a minimum clip duration of 3.5 seconds, ensuring each clip is long enough to encompass a full action. Simultaneously, to avoid excessively long clips, which can complicate analysis and may contain extraneous information, we capped the maximum duration at 10 seconds. For videos exceeding this length, only the most relevant 10-second segment, typically featuring the most significant action, was selected. This methodology results in a dataset that is both comprehensive and tailored for efficient action recognition analysis.
The following table details the distribution of video durations within our dataset, post these adjustments:
Duration Range (seconds) Number of Videos
3.5-5 30103
5-10 81356
10+ 172123
Table 6: Distribution of video durations in the TikTokActions dataset. The data reflects the adherence to the minimum and maximum duration thresholds for clip selection.