### v3

# Advancing Human Action Recognition with Foundation Models trained on Unlabeled Public Videos

Abstract
The increasing variety and quantity of tagged multimedia content on a variety of online platforms offers a unique opportunity to advance the field of human action recognition. In this study, we utilize 283,582 unique, unlabeled TikTok video clips, categorized into 386 hashtags, to train a domain-specific foundation model for action recognition. We employ VideoMAEv2, an advanced model integrating Masked Autoencoders (MAE) with Vision Transformers (ViT), pre-trained on this diverse collection of unstructured videos. Our model, fine-tuned on established action recognition benchmarks such as UCF101 and HMDB51, achieves results comparable with the state-of-the-art: 97.5% on UCF101, 83.9% on HMDB51, 81.1% on Kinetics-400, and 69.2% on Something-Something V2 using the ViT-base backbone. These results highlight the potential of using TikTok videos as a valuable source of diverse and dynamic content for training foundation models. Our investigation confirms that while initial increases in pre-training data volume significantly enhance model performance, the gains diminish as the dataset size continues to expand. Our findings identify two critical lessons in self-supervised learning for computer vision: (1) additional pre-training data yields diminishing benefits and (2) these diminishing benefits can be observed at relatively small dataset sizes on the order of thousands of images.
1 Introduction
Action recognition is a fundamental computer vision task that involves identifying and classifying human actions in video sequences, with significant applications in areas such as security, interactive gaming, and healthcare. In the era of social media platforms, TikTok has emerged as a significant source of human action videos. We utilized a collection of 283,582 TikTok video clips, encompassing a wide range of human activities, to train a foundation model tailored to human action recognition. This internally curated collection aims to represent various human actions relevant to real-world scenarios and cultural contexts. Importantly, it includes well-defined but non-standard human actions, addressing gaps in existing action recognition datasets and diversifying the range of human activities recognized by current models.
Unlike Kinetics [4], which also includes user-generated content from YouTube, our collection exclusively features TikTok videos, known for their dynamic and spontaneous nature. This distinction offers unique training advantages, as such videos often exhibit quick transitions, novel perspectives, and creative interactions—elements less prevalent in YouTube’s often longer-form content. These characteristics provide robust challenges for training models, enhancing their ability to interpret a wide range of human actions in diverse and unpredictable environments.
We conducted a series of experiments to assess the potential of videos collected in this manner for pre-training computer vision foundation models for action-related tasks. We employed the well-established, state-of-the-art model VideoMAE V2, using it with ViT-base backbones[9]. The ViT-base was used to provide a baseline comparison with other models using similarly sized backbones. Subsequent fine-tuning on well-known video action recognition datasets like UCF101[7], HMDB51[5], Kinetics[4], and Something-Something V2[3] demonstrated valuable insights: the performance achieved with both backbones was comparable to, or in some cases even better than, models trained exclusively on these conventional datasets or even those pre-trained on much larger datasets, where performance increases are often marginal.
We also explored the relationship between the quantity of pre-training videos and the performance on downstream tasks. By sampling subsets of our collection, starting with 1,000 videos and increasing in increments of 1,000 up to 6,000 videos, and pre-training VideoMAE V2[9] on these subsets, we observed that performance remains robust as long as the number of pre-training videos stays above a certain threshold, highlighting that massive datasets may not always be necessary for effective pre-training. This exploration not only emphasizes the efficiency of using online unlabeled videos as a pre-training source but also challenges the conventional reliance on extensive video datasets for achieving high model performance. Before deployment of this approach in the real world, we encourage researchers to think through the ethics of data scraping from publicly available sources to train deep learning models.
2 Related Work
2.1 Human Action Recognition Datasets
The evolution of action recognition research has been strongly influenced by the availability of benchmark datasets. The UCF101[7] dataset, with its diverse collection of human actions, has been a cornerstone for early research in this domain, guiding and benchmarking numerous computer vision innovations applied to action recognition. Similarly, the HMDB-51[5] dataset, despite its limited size, has been pivotal in providing a varied set of human actions captured under realistic conditions.
The Kinetics Human Action Video dataset[4], introduced in 2017, marked a significant milestone in this field. With its vast collection of YouTube-sourced videos encompassing 400 distinct human action classes, Kinetics has become an indispensable resource for researchers, setting a higher benchmark and promoting advanced research in action recognition.
Another noteworthy dataset is the Something-Something[3] collection. Its unique focus on human-object interactions captured in real-world scenarios has offered researchers a different perspective, emphasizing the complexity and diversity of everyday actions.
These datasets, along with several others, have collectively laid the foundation for the current advancements in action recognition, facilitating the development and evaluation of state-of-the-art models and techniques. Recognizing the rapid evolution of content on platforms like TikTok and the emergence of diverse, contemporary actions and scenarios, we are motivated to develop a comprehensive dataset highlighting the latest trends and user behaviors specific to new short-video social media platforms. Our aim was to provide the community with a dataset with well-defined yet non-standard human actions that complement common actions already represented in current action recognition datasets, thus adding variety to the actions that action recognition models are currently evaluated for.
2.2 Computer Vision Foundation Models
Foundation models have become essential tools in the field of computer vision. The advent of models like Florence[10] showcases the versatility and scope of these models. Specifically, Florence encapsulates the progression of foundation models, adeptly offering transitions from broad scene outlines to detailed object-centric views and from static imagery to dynamic video sequences. Additionally, its effectiveness spans various modalities, from RGB imagery to textual data, establishing new standards in tasks like classification, object detection, retrieval, and particularly in action recognition.
Another significant model in this domain is VideoMAE V2[9]. Introduced at CVPR 2023, this self-supervised pre-training methodology stands out for its ability to train on datasets exceeding 1 million videos. With its innovative dual masking strategy and progressive training approach, VideoMAE V2 sets new standards in action recognition. Empirical evaluations demonstrate its superior performance across multiple benchmarks, emphasizing its effectiveness in real-world scenarios. We anticipate that our extensive TikTok dataset will serve as a valuable resource for the continued development and refinement of computer vision foundation models, contributing to the advancement of this rapidly evolving field.
3 Dataset Construction
3.1 Video Collection Overview
To support our experimentation exploring the utility of unlabeled public data for pre-training foundation models for human action recognition, we curated an extensive collection of TikTok videos. This collection, consisting of over 280,000 video clips from 386 hashtags, captures a wide spectrum of human actions across various scenarios. The videos span diverse categories, from dance movements and fitness exercises to hand gestures and daily activities. Unlike traditional datasets like Kinetics, our collection uniquely incorporates elements of pop culture, memes, and rapid trend evolution, providing dynamic content that significantly boosts the model’s ability to understand and generalize from real-world, culturally relevant scenarios.
3.2 Video Selection Methodology
To curate a diverse and relevant set of videos from TikTok, we initially examined several well-known action recognition datasets, including UCF101[7], HMDB51[5], and Google DeepMind Kinetics[4]. These datasets helped us to understand the broad categories of human actions typically represented in research. We identified four primary categories: dance, sports, fitness, and kinetics.
Building upon these categories, we explored TikTok to identify relevant hashtags that capture a wide array of human actions within these domains. Our selection was strategically guided by six principles designed to ensure a comprehensive and culturally relevant collection: generalization, specification, noun substitution, noun-verb order, hashtag consolidation, and social/cultural relevance. These principles helped ensure that the videos not only cover a broad spectrum of actions but also include dynamic and culturally significant content such as memes, pop culture trends, and unique challenges—elements that are distinctly prevalent on TikTok.
The social/cultural relevance was particularly important, allowing us to incorporate an additional category that includes a wide variety of actions originating from distinct social and cultural contexts, such as viral dance challenges and fitness trends. Popularity and engagement levels also played a crucial role in our selection process; we set a minimum threshold of 5 million views per hashtag to ensure that the content is not only relevant but also resonates widely with users, reflecting substantial public engagement and visibility. This approach allows us to harness rich, culturally nuanced content that can aid in training models to better understand and interpret human actions in real-world contexts.
3.2.1 Generalization/Using the umbrella terms
When comparing the hashtags we generated with those used by TikTok users, we observed that hashtags related to specific actions can often be represented in a very broad or inclusive manner on the platform. For instance, the hashtag #basketball can serve as an umbrella term encompassing a wide array of hashtags, such as #playingbasketball, #dunking, #dribbling.
3.2.2 Specification
There are situations where specificity is crucial. For instance, within the broader classification of Ballroom Dance, distinct dance styles such as Waltz, Quickstep, and Tango exist. Instead of solely using the hashtag #ballroomdance, we include specific hashtags, including #chacha, #salsa, #tango, and #waltz, to ensure our datasets encompass pertinent human actions inherent to distinct dance styles.
3.2.3 Noun Substitution
While gathering short TikTok videos, we observed that a more effective approach for identifying the desired content is to employ a single noun instead of an entire verb phrase. For instance, the use of the hashtag #piano suffices to compile a more extensive collection(see Table  1) of short videos instead of using #pianoplaying.
Hashtags Number of Views
#piano 44.4 Billion
#playingpiano 29.4 Million
Table 1: Example of the expansion in data availability when removing nouns from the hashtag construction process.
Figure 1: The hashtag selection and modification process.
3.2.4 Noun-Verb Order
We noticed a preference among users for using noun phrases over verb phrases when describing their activities. For example, TikTok users frequently use hashtags like #shoerepair instead of verb phrases like #repairingshoes.
3.2.5 Hashtag Consolidation
When composing the hashtag list, we streamlined similar hashtags into a single one if they denote the same type of human action, prioritizing those with higher views on TikTok. For example, we retain #cuttingcakes and remove similar hashtags like #cutfruit, #applecutting, and #breadcutting.
3.2.6 Social/cultural Considerations
Culturally and socially relevant hashtags were included when crafting the hashtag list. For instance, the ”flower challenge” originated as a dance challenge on TikTok initiated by a K-pop star, and TikTok users subsequently adopted the hashtag #flowerchallenge when sharing their dance videos.
In Figure 1, we show an example flowchart of the hashtag selection and modification process. Our final and complete list of hashtags along with video numbers is in Appendix A.
3.3 Video Collection and Processing
We curated a comprehensive video collection using the official TikTok research API, focusing on approximately 900 videos per hashtag from a set of hashtags identified through our selection process. Each video and its metadata were processed to facilitate model training, using tools such as PySceneDetect and YOLO v8[6]. This processing resulted in a JSON file for each video, which included metadata capturing attributes like title, duration, and scene distribution, as analyzed by PySceneDetect.
3.3.1 PySceneDetect and YOLOv8 Processing
Our processing approach began with PySceneDetect, used to analyze the scene composition of each video. This tool helped identify distinct scenes within the videos, enabling us to focus on the longest scene clips, hypothesized to have the highest likelihood of containing significant human action.
Once the longest scenes were identified, the YOLOv8[6] model was employed to detect human figures in these segments. The primary goal of this processing was to ensure accurate identification of videos containing human subjects, vital for the training of our models.
During the frame extraction phase, we used the ffprobe tool to determine the total frame count of each video clip. Due to the large number of frames in many videos, we applied subsampling for efficiency. Typically, a fixed number of frames (usually 10) were extracted at regular intervals throughout the video.
The extracted frames were then processed using the YOLOv8[6] model to detect human presence. Frames that included at least one ”person” label were counted towards the total number of frames containing humans for that particular video.
Post-processing included applying a threshold for human presence; videos needed to exceed a pre-defined frame threshold with human detections to qualify. Videos that did not meet this threshold were subject to further review. This systematic approach helped compile aggregated statistics on the human presence in our video collection, ensuring the quality of content used for model training. The entire process, from video selection to frame extraction and human detection, was documented to maintain transparency and reproducibility. Figure  2 illustrates the process of clip extraction from the raw videos sourced from TikTok.
Figure 2: The clip extraction process from raw videos mined on TikTok.
4 Performance on Benchmarks
In the domain of human action recognition, benchmarking performance is essential to understand the efficacy of a model. This section evaluates the performance of our model, initially pre-trained on the entire collection of 283,582 unlabeled TikTok video clips using VideoMAE V2[9], and subsequently fine-tuned on Kinetics-400 (K400). Following this, the model underwent further fine-tuning on UCF-101[7] and HMDB51[5], adopting a transfer learning approach akin to the experimental setups used in VideoMAE V2 studies. Additionally, we extended our evaluation to include the Something-Something V2[3] dataset, assessing the model’s adaptability and performance across varied action recognition contexts.
4.1 Pre-training on VideoMAE V2
VideoMAE V2[9] is a state-of-the-art masked autoencoding framework, notable for its flexibility in backbone selection, ranging from base to billion-level configurations. We opted for the ViT-base backbone, which, while less computationally intense than its billion-level counterparts, significantly enhances performance across various video-related tasks, including video action recognition. The generalization capability of VideoMAE V2 pre-trained ViTs[1] as video foundation models has been demonstrated across multiple downstream tasks, highlighting its robustness and adaptability.
Our training was conducted on 8 NVIDIA A100 40GB GPUs, utilizing the full set of 283,582 unlabeled TikTok videos. The pre-training on the ViT-base backbone lasted for 20 days, providing a comprehensive base for developing a robust model capable of understanding a wide array of human actions.
4.2 Fine-Tuning on Established Datasets
After initial pre-training, the model underwent two distinct paths of fine-tuning. The first path involved direct fine-tuning on the Kinetics-400 and Something-Something V2 datasets to assess immediate applicability to broad action recognition tasks. The results of this direct fine-tuning are detailed in Table 2. The second path involved using the model fine-tuned on Kinetics-400 as an intermediary for transfer learning on UCF-101 and HMDB51, leveraging robust pre-training on TikTok videos. The outcomes of this transfer learning process are presented in Table 3.
Dataset Top-1 (%) Top-5 (%)
Kinetics-400[4] 81.1 94.8
Something-Something V2[3] 69.2 91.6
Table 2: Direct Fine-Tuning Results With TikTok Videos Pre-training
Dataset Top-1 (%) Top-5 (%)
HMDB-51[5] 83.9 96.7
UCF-101[7] 97.5 99.6
Table 3: Transfer Learning Results From K400 Fine-Tuning Through TikTok Pre-training
The fine-tuning results highlight the significant potential of using large, unlabeled video collections from platforms like TikTok for advancing human action recognition tasks. This approach capitalizes on the rich, varied nature of content available on social media, which, when used for pre-training, enhances the model’s ability to generalize across different datasets and recognition scenarios. The robust performance across both direct and transfer learning paths underscores the value of TikTok videos in training models to tackle complex action recognition challenges.
5 Extrapolating the Required Pre-Training Dataset Size
In Figure LABEL:fig:accplot, we find a monotone increasing accuracy score versus data volume. In the spirit of Floares et al. [2], we attempt to extrapolate the number of preprocessed videos for a desired accuracy by curve fitting to the number of data points trained so far. Because the accuracy maxes out at 100%, we choose to fit curves that are monotonically increasing with a horizontal asymptote at $11$. This makes any CDF of a random variable a viable choice for curve fitting. By iterating over the random variables in Scipy [8], we choose the top 10 distributions that fit the accuracy curve. Table 4 displays the number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
We validated these projections by pre-training a model with 9157 data points and fine-tuning the model using the same methods applied in the initial projections. The outcome of this validation was notably consistent with our predictive model. The model achieved an accuracy of 94.93%, closely aligning with the expected performance as indicated by the inverse CDF projections for this dataset size.
This validation not only reinforces the reliability of our statistical model but also highlights the practicality of using such predictive techniques in determining dataset size requirements. The fact that a dataset size of just 9157 points could yield an accuracy so close to 95% – a benchmark often targeted in action recognition tasks – is a testament to the efficiency of our approach. These results suggest that the requirement for massive datasets in the pre-training phase may be reconsidered, especially when computational resources are limited. This could mark a significant shift in how datasets are compiled and utilized in machine learning, particularly in fields where data acquisition is a challenge.
Table 4: Number of pre-training data points required to achieve high accuracy levels (between 95% and 99%) according to the inverse CDF for common statistical models.
Model Desired Accuracy
95% 96% 97% 98% 99%
t 9157 12388 18323 31892 82574
Burr XII 9090 12209 17857 30511 76240
InvGamma 9072 12166 17750 30204 74859
BetaPrime 9072 12165 17748 30200 74843
F 9072 12166 17749 30203 74855
GenPareto 9072 12164 17744 30184 74759
Lomax 9072 12164 17744 30184 74759
LogLaplace 9072 12164 17744 30184 74760
Pareto 9072 12164 17743 30183 74755
InvWeibull 9043 12101 17596 29772 72915
6 Discussion
The curation of TikTok videos for enhancing human action recognition models provides a unique contribution to the field, especially by incorporating a variety of distinct yet non-traditional human actions. These actions, sourced from the TikTok platform, were carefully selected to complement the more conventional actions typically found in established datasets like UCF101[7] and HMDB51[5]. This approach not only diversifies the types of human actions available for model training but also qualitatively enriches the dataset, thereby challenging and extending the adaptability of action recognition algorithms.
Integrating these unique and diverse actions from TikTok, we have expanded the range of actions that machine learning models can recognize, addressing the dynamic nature of human actions in digital media where new and unconventional actions frequently emerge. This strategy highlights the value of leveraging culturally rich and dynamic user-generated content as a training resource in an era dominated by digital media.
Furthermore, our findings challenge the conventional wisdom that larger datasets always yield better pre-training results. The competitive performance of our models, initially pre-trained on a curated subset of TikTok videos and subsequently fine-tuned on Kinetics-400[4] (K400) and Something-Something V2(Sth-Sth)[3], suggests that a well-curated, smaller dataset can sometimes outperform a larger, more generic one. This observation encourages a shift towards more efficient dataset compilation and model training paradigms, emphasizing data quality over quantity.
Looking ahead, exploring the potential of using the full breadth of the TikTok video collection for pre-training could further enhance model performance and generalization across various tasks. Additionally, integrating these videos with other datasets, such as Kinetics-400 and Something-Something, could provide deeper insights into the unique characteristics and potential synergies between different types of video content.
An emerging area of interest is the application of weekly self-supervised learning techniques, which involve regular, unsupervised updates to the model based on new data inflows. Such techniques could significantly boost the adaptability and accuracy of models trained with dynamic, culturally relevant content from platforms like TikTok.
7 Ethics
This research adheres to strict ethical guidelines regarding the use of online video data, specifically videos collected from TikTok. All videos were legally accessed via the TikTok Research API, and only those videos where the user granted download permission were included. Our data collection process is in full compliance with the TikTok Research API Terms of Service. As an additional layer of security, we delete the data after training.
We place a high priority on privacy and copyright issues. To protect individual privacy, we have implemented face obfuscation techniques in the video clips included in our paper. Moreover, in line with the terms of service, we do not distribute any TikTok user content directly. Instead, we focus on training a foundation model, which we only use to identify action patterns, not personal biometric features.
Ethical Considerations in Online Video Research: The use of online video data poses unique challenges and ethical considerations, particularly regarding privacy, informed consent, and the potential risks associated with the use of such data. Our study emphasizes the importance of adhering to ethical principles that safeguard participant interests. We have made significant efforts to minimize potential harm by anonymizing the data and not sharing the scraped dataset. Although techniques like pixelating faces can be circumvented and are not foolproof, they represent our commitment to reducing identifiable information.
Furthermore, we advocate for the inclusion of clearer guidelines in the terms of service on video platforms, which would allow users to explicitly exclude their videos from research use. This approach does not replace the need for informed consent but is a step towards enhancing ethical practices in online video research. We aim to foster a transparent and informed discussion about the use of online video data in computational research.
Acknowledgements
This work was made possible by a multitude of supporters and funding sources. We thank the National Science Foundation for their generous support through the CyberTraining grant (Award No. 2118222) under the HI-DSI project. The technical assistance and advanced computing resources provided by the University of Hawaii Information Technology Services–Cyberinfrastructure, partially funded by NSF CC* awards #2201428 and #2232862 (Koa), have been invaluable. The development and acceleration of cloud computing through JetStream2 have significantly contributed to our research; we acknowledge David Y. Hancock, Jeremy Fischer, John Michael Lowe, Winona Snapp-Childs, Marlon Pierce, Suresh Marru, J. Eric Coulter, Matthew Vaughn, Brian Beck, Nirav Merchant, Edwin Skidmore, and Gwen Jacobs for this computational infrastructure (DOI: 10.1145/3437359.3465565). The AWS Cloud Credit for Research program has also played a crucial role in providing computational resources for our work. We are grateful for the funding provided by the National Institutes of Health under the NIH Director’s New Innovator Award (award DP2-EB035858). We used ChatGPT to edit the grammar of our manuscript and to re-phrase sentences that were originally worded unclearly. However, all contents in this manuscript are original ideas and analyses conducted by the authors.
Appendix A Appendix
In the appendix, we will present additional insights into our dataset, specifically focusing on visual and statistical aspects. This includes select screenshots of our video clips, which exemplify the diversity and context of the actions captured within our dataset. Alongside these visuals, we will include the figure, which provides the number distributions of videos for each hashtag. Furthermore, we will present the distribution of video durations, offering a comprehensive view of the temporal characteristics of our clips. These elements combined will give a more complete picture of the dataset’s richness.
Appendix B Sample Clips
In Figure 3, we present the extensive diversity of the TikTokActions dataset, which captures a wide range of human activities tagged across multiple domains such as fitness, sports, and dance. The collection features a variety of video types that include full-body human actions, clips focusing on hand movements, and scenarios involving multiple individuals. The presence of such varied content provides a rich learning ground for self-supervised learning models to extract and learn a comprehensive set of features from human actions, gestures, and interactions, aiding in the enhancement of action recognition capabilities.
Figure 3: Example classes from the TikTokActions dataset.
Appendix C Dataset Composition
Our dataset offers a diverse array of video counts across different hashtags. Figure 4 illustrates the distribution of videos across various hashtags, indicating the prevalence of certain actions within our dataset. On average, each hashtag is associated with approximately 735 videos. The ’sax’ category stands out with the maximum count of 938 videos, while the ’mopping’ category has the minimum, with 325 videos.
Figure 4: Bar chart showing the number of videos per hashtag in our dataset, demonstrating the frequency and distribution of action categories.
Conversely, Figure 5 displays a pie chart that categorizes the dataset into four video count ranges: 200-400, 400-600, 600-800, and 800+ videos per category. It is noteworthy that a significant portion, 53.4%, of the hashtags have a video count that falls within the 600-800 range.
Figure 5: Video counts in different ranges, highlighting the dataset’s variety and the volume of data available for model training.
Appendix D Dataset Duration
In the composition of the TikTokActions dataset, a key focus was on ensuring that each video clip is sufficiently long to capture a complete action, while also maintaining a manageable length for analysis. To this end, we established a minimum clip duration of 3.5 seconds, ensuring each clip is long enough to encompass a full action. Simultaneously, to avoid excessively long clips, which can complicate analysis and may contain extraneous information, we capped the maximum duration at 10 seconds. For videos exceeding this length, only the most relevant 10-second segment, typically featuring the most significant action, was selected. This methodology results in a dataset that is both comprehensive and tailored for efficient action recognition analysis.
The following table details the distribution of video durations within our dataset, post these adjustments:
Duration Range (seconds) Number of Videos
3.5-5 30103
5-10 81356
10+ 172123
Table 5: Distribution of video durations in the TikTokActions dataset. The data reflects the adherence to the minimum and maximum duration thresholds for clip selection.